OpenStudy (anonymous):

angle MNP~angle QRS, and the ratio of angle MNP TO QRS is 5:2.if MN=42meters,what is the QR?

OpenStudy (anonymous):

angle mnp / angle qrs = MN / QR = 5/2 = 42/QR = QR = 21/5 I THINK !!!

OpenStudy (anonymous):

thankQ

Directrix (directrix):

Triangle MNP ~ Triangle QRS With scale factor 5 to 2 MN/QR = 5/2 42/QR = 5/2 5 QR = 84 QR = 84/5

OpenStudy (anonymous):

thankQ Directrix

Directrix (directrix):

Glad to help. Hope that I did not add too much in the translation.

OpenStudy (anonymous):

ya i am glad to get help from u,any ways Its okay.

OpenStudy (anonymous):

OOOOOOOOOOOOOOOOOOOOOOOOOOH, YES Directrix , it is 84/5 , ma answer came 21 / 5 because i did little bit mistake in cross multiplication it is : 5/2=42/QR and here we have 2 multiply 5 with 42 , i did a mistake here , hope u both got it !

OpenStudy (anonymous):

anyways i am glad to get help from u mugeesULkaisar anyways thankyyou

OpenStudy (anonymous):

welcome !